OpenStudy (anonymous):

3. What is the value of x in the regular polygon below?

OpenStudy (anonymous):

40 120 60 150 ?

OpenStudy (anonymous):

what is the sum of all angles in a triangle

OpenStudy (anonymous):

no idea?

OpenStudy (anonymous):

(n-2)*180 .... n is the number of sides (in this case 6) (6-2)*180 = 4*180 = 720 total 720/6 = 120 for each angle so 3x = 120 ....divide by 3 x =40

OpenStudy (anonymous):

Thanks :) Lol